OpenStudy (riley):

The RideEm Bicycles factory can produce 100 bicycles in a day at a total cost of $10,400 and it can produce 120 bicycles in a day at a total cost of $11,200. What are the company's daily fixed costs? What is the marginal cost per bicycle?

OpenStudy (anonymous):

same as saying "find the equation for the line between (100,10400) and (120,11200)

OpenStudy (anonymous):

slope is the "marginal cost" and the y intercept is the fixed cost

OpenStudy (riley):

O_O . . . I have the sudden urge to use the slope formula . . .

OpenStudy (anonymous):

good plan

OpenStudy (riley):

Thank you satellite<3

OpenStudy (riley):

I just did three of these a min ago, but it wasn't a "word problem," which totally tripped me up. I think I got this now. xD

OpenStudy (anonymous):

and it actually makes total sense because an extra 20 bikes cost an extra $800 so you get a price per bike of $40. notice that this is the same as saying \[\frac{11200-10400}{120-100}=40\]

OpenStudy (riley):

Yeah I just realized it when I plugged it all into the slope formula. >_< Then I just go back and plug the slope into the equation y=mx+b, for m. Then use one of the number pairs to find b right?

OpenStudy (riley):

Working it out, I think I got it now.

OpenStudy (riley):

Okay so: Marginal would = 40 Fixed cost = 6400

OpenStudy (anonymous):

yup and that makes sense too. because you know each bike cost $40 to 100 bikes cost $4000 but the actual cost was $10400 to you know the fixed cost was $10400-$4000= = $6400

OpenStudy (riley):

ahhhhhh I'm seeing the big picture now xD

OpenStudy (anonymous):

i guess what i am saying is that the "word problem" makes sense out of finding the equation of the line. you might could have done this without even thinking "find the slope, find the y intercept"

OpenStudy (riley):

Yeah, it's like show me a problem like this in equation form and I can find it easy peezy. But throw words in there and turn it into a world problem . . . and it totally scrambles my mind. >_< I think I need to learn to look at word problems differently or something, like see the problem within the problem . .

OpenStudy (riley):

thank you again for the help, it makes complete sense after you explained it like that